OpenStudy (keeli_grace):

WILL FAN AND MEDAL!!! :) What is the slope of the line that passes through the points (3, 1) and (–2, 5)? A. -5/4 B. 5/4 C. -4/5 D. 4/5

OpenStudy (anonymous):

c bc you do 5-1/-2-3 which = -4/5

OpenStudy (anonymous):

\[\large slope=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

keeli can you help me with mine or orange master

OpenStudy (anonymous):

\[\large slope=\frac{ 5-1 }{ (-2)-3 }\]\[5-1=4\]\[(-2)-3=-5\]\[\large slope=\frac{ 4 }{ -5 }=C\]

OpenStudy (keeli_grace):

Thank you!!! :)